A body is projected with a velocity u. It passes through a certain point above the ground after t1 sec . The time after which the body passes through the same point during the return of journey is : ?

A body  is projected with a velocity u. It passes through a certain point above the ground after t1 sec . The time after which the body passes through the same point during the return of journey is : ?

Grade:12th pass

3 Answers

Arun
4 years ago
Let the starting point be A, point crossed by the ball be B and final point from where the ball returns be C. 
For AC, v=0, u=u, and total time= t
0= u-gt
Or t= u/g ---eq.1
Now from A to B, time = t1 & from B to C, t= t2.
t = t1+t2 or t2 = t - t1.
But the ball comes back to the point B from C,
Implies that t2= 2(t- t1)
= 2(u/g - t1) is the final answer.
 
 
 
Khimraj
4 years ago
A body is projected upwards with a velocity u. 
It passes through a certain point above the ground after T1 second the time after which the body pass through the same point during the return journey is 2u/g - t1. 
t1 + t2= 2u / g 
So t2 =  2(u/g - t1)
aswanth nayak
4 years ago
Dear Student,
 
A body is projected upwards with a velocity u. 
It passes through a certain point above the ground after T1 second the time after which the body pass through the same point during the return journey is 2u/g - t1. 
t1 + t2= 2u / g 
So t2 =  2(u/g - t1)
 
Hope it helps you
Regards